4) Find the vector function that represents the curve of intersection for the following surfaces (assume y > 0): x2 + z2 = 9We will all rate if correct
Answer #1

a²+ z²=g and ac²t}+42²=25 2²+z²=9 x=3cost 3 sint lost < 2t z = a?+y+42= 25 (+23H+ y +327=25 9+ y +37= 25 y²+ 32²=16 y = 16-32